1/2x-1/4y=-1
2(y-1)-2(1-x)=y+4

Answer by IBRAHIM92(5)

1/2x-1/4y=-1
2(y-1)-2(1-x)=y+4
Multiply the first equation x4:
2x-y=-4
Remove the brackets in the second equation:
2y-2-2+2x=y+4
y+2x=8
now the equations :
2x-y=-4
y+2x=8
Add the two equations :
4x =4 , Hence x=1
then :
y =6